# Vulnerability of LLMs to Vertically Aligned Text Manipulations

### v1

@@ 4.5 Experiment Results
        Model         SST-2         CoLA         QNLI         Rotten Tomatoes         Jigsaw Toxicity         Closed-Source Models         GPT-3.5  w/ CoT         61.00 $(\downarrow\textbf{4.00})$         50.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         59.00 $(\downarrow\textbf{10.00})$         53.00 $(\downarrow\textbf{4.00})$         62.00 $(\textbf{0.00})$         GPT-4  w/ CoT         66.00 $(\downarrow\textbf{1.00})$         51.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         68.00 $(\downarrow\textbf{3.00})$         60.00 $(\downarrow\textbf{4.00})$         56.00 $(\downarrow\textbf{2.00})$         GPT-4o-mini  w/ CoT         68.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         50.00 $(\textbf{0.00})$         76.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$         64.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         56.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         GPT-4  w/ CoT         71.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         52.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{5.00}})$         74.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{4.00}})$         66.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$         66.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{6.00}})$         Open-Source Models         Llama3-8B  w/ CoT         59.00 $(\downarrow\textbf{2.00})$         49.00 $(\downarrow\textbf{1.00})$         64.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         53.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{11.00}})$         55.00 $(\downarrow\textbf{3.00})$         Llama3-70B  w/ CoT         66.00 $(\downarrow\textbf{1.00})$         46.00 $(\downarrow\textbf{2.00})$         59.00 $(\downarrow\textbf{4.00})$         55.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{9.00}})$         62.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{4.00}})$         Llama3.1-8B  w/ CoT         53.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         51.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         62.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         56.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         63.00 $(\downarrow\textbf{1.00})$         Llama3.1-70B  w/ CoT         68.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         47.00 $(\downarrow\textbf{3.00})$         65.00 $(\downarrow\textbf{1.00})$         62.00 $(\downarrow\textbf{1.00})$         65.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         Gemma2-9B  w/ CoT         60.00 $(\textbf{0.00})$         49.00 $(\downarrow\textbf{2.00})$         64.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{4.00}})$         61.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         52.00 $(\downarrow\textbf{1.00})$         Gemma2-27B  w/ CoT         61.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         51.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$         64.00 $(\textbf{0.00})$         57.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$         53.00 $(\downarrow\textbf{2.00})$         Qwen1.5-72B  w/ CoT         63.00 $(\textbf{0.00})$         51.00 $(\downarrow\textbf{1.00})$         63.00 $(\downarrow\textbf{3.00})$         59.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         60.00 $(\downarrow\textbf{3.00})$         Qwen2-72B  w/ CoT         62.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         52.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         64.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$         60.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$         61.00 $(\uparrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$
@@
Table 4: Accuracy scores after applying CoT, with parentheses indicating the change in scores (compared to the non-CoT scores in Table 1). Red indicates an increase in accuracy, whereas black denotes either no change or a decline in accuracy. All other experimental settings remain consistent with those in Table 1.
@@
To ensure comprehensive experiments, we also test three high-performing open-source LLM series: Llama, Gemma, and Qwen. As shown in the lower part of Table 1, these open-source models also display considerable vulnerability to vertically formatted text inputs. This suggests that current LLMs are not trained with consideration for the impact of vertical text inputs, resulting in a high degree of vulnerability, which may pose certain risks in some areas.
@@ 5.1 Failure of CoT
In the above experiments, we have the LLMs directly output classification results for two types of text inputs. This method somewhat bypasses the model’s reasoning process, which may result in inaccuracies. To address this, we consider using a Chain of Thought (CoT) approach Wei et al. (2023) to determine if prompting the model to output reasoning steps can mitigate its vulnerability to vertically formatted text inputs. We incorporate "Think step by step" into the prompt to encourage the LLMs to provide a step-by-step reasoning process, aiming to help the model better understand the intricacies of vertically formatted text inputs.
@@ 6.1 Disordered Tokens
Due to the natural left-to-right and top-to-bottom writing order of text, the tokenizer used by LLMs encodes input text in this sequence. For instance, the tokenizer for the Llama3-8B model represents the word "vertical" as a single token with an ID of 15651 when input horizontally. However, when input vertically, it converts into a token sequence of length 15 due to the multiple spaces and line breaks. This causes the LLM to lose its understanding of the complete word.
@@ 6.2 Lost Attention
In summary, the tokenization conventions of LLMs (processing text from left to right and top to bottom) and the lack of relevant pre-training data have impaired their ability to understand vertical text. In contrast, people can easily comprehend words presented in a vertical format. Consequently, the cognitive disparity between humans and LLMs leads to the models’ vulnerability to vertical text input, which could pose potential threats that humans might exploit.

### v3

@@ 4.4 Experiment Results
      Model       SST-2       CoLA       QNLI       Rotten Tomatoes       Jigsaw Toxicity       Closed-Source Models       GPT-3.5  w/ CoT       61.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{4.00}})$       50.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       59.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{10.00}})$       53.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{4.00}})$       62.00 $(\textbf{0.00})$       GPT-4  w/ CoT       66.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$       51.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       68.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$       60.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{4.00}})$       56.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$       GPT-4o-mini  w/ CoT       68.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       50.00 $(\textbf{0.00})$       76.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{1.00}})$       64.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       56.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       GPT-4o  w/ CoT       71.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       52.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{5.00}})$       74.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{4.00}})$       66.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{1.00}})$       66.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{6.00}})$       Open-Source Models       Llama3-8B  w/ CoT       59.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$       49.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$       64.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       53.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{11.00}})$       55.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$       Llama3-70B  w/ CoT       66.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$       46.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$       59.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{4.00}})$       55.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{9.00}})$       62.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{4.00}})$       Llama3.1-8B  w/ CoT       53.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       51.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       62.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       56.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       63.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$       Llama3.1-70B  w/ CoT       68.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       47.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$       65.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$       62.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$       65.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       Gemma2-9B  w/ CoT       60.00 $(\textbf{0.00})$       49.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$       64.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{4.00}})$       61.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       52.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$       Gemma2-27B  w/ CoT       61.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       51.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{1.00}})$       64.00 $(\textbf{0.00})$       57.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{3.00}})$       53.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{2.00}})$       Qwen1.5-72B  w/ CoT       63.00 $(\textbf{0.00})$       51.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{1.00}})$       63.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$       59.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       60.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{3.00}})$       Qwen2-72B  w/ CoT       62.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       52.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       64.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$       60.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{1.00}})$       61.00 $(\uparrow{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\textbf{2.00}})$
@@
Table 4: Accuracy scores after applying CoT, with parentheses indicating the change in scores (compared to the non-CoT scores in Table 1). Dark green indicates an increase in accuracy, red denotes a decline, and black means no change. All other experimental settings remain consistent with those in Table 1.
@@
To ensure comprehensive experiments, we also test three high-performing open-source LLM series: Llama, Gemma, and Qwen. As shown in the lower part of Table 1, these open-source models also display considerable vulnerability to vertically formatted text inputs. Moreover, on four experimental datasets, Llama3.1-8B shows a greater accuracy drop than Llama3-8B, despite being a stronger language model. This suggests that current LLMs are not trained with consideration for the impact of vertical text inputs, resulting in a high degree of vulnerability, which may pose certain risks in some areas.
@@ 5.1 Failure of CoT
In the above experiments, we have the LLMs directly output classification results for two types of text inputs. This method somewhat bypasses the model’s reasoning process, which may result in inaccuracies. To address this, we consider using a Chain of Thought (CoT) approach Wei et al. (2023) to determine if prompting the model to output reasoning steps can mitigate its vulnerability to vertically formatted text inputs. We incorporate "think step by step" into the prompt to encourage the LLMs to provide a step-by-step reasoning process, aiming to help the model better understand the intricacies of vertically formatted text inputs Kojima et al. (2023).
@@ 6.1 Disordered Tokens
Due to the natural left-to-right and top-to-bottom writing order of text, the tokenizer used by LLMs encodes input text in this sequence. For instance, the tokenizer for the Llama3.1-8B model represents the word "vertical" as a single token when input horizontally. However, when input vertically, it converts into a token sequence of length 15 due to the spaces and line breaks. This causes the LLM to lose its understanding of the complete word.
@@ 6.2 Lost Attention
In summary, the tokenization conventions of LLMs and the lack of relevant pre-training data have impaired their ability to understand vertical text. In contrast, people can easily comprehend words presented in a vertical format. Consequently, the cognitive disparity between humans and LLMs leads to the models’ vulnerability to vertical text input, which could pose potential threats that humans might exploit.
@@ 7 Conclusion
Acknowledgments
This work is supported by NSF #2331966, DARPA ANSR program FA8750-23-2-0004, an Amazon AGI Research Award, and University of California, Merced.
Appendix A Large Language Models
  • GPT-3.5: A robust large language model developed by OpenAI, capable of generating text based on instructions, and highly effective across diverse natural language processing tasks.
  • GPT-4: An advanced multi-modal language model from OpenAI that accepts both image and text inputs for text generation, achieving near-human performance on various benchmarks.
  • GPT-4o-mini: A cost-efficient multimodal model, released by OpenAI on July 18, 2024, is a distilled version of GPT-4o, offering low latency and cost while supporting a wide range of tasks.
  • GPT-4o: OpenAI’s latest multimodal AI model, offering enhanced reasoning, generation, and understanding capabilities across text, image, and speech with faster and more efficient responses.
  • Llama: Meta’s open-weight AI models, designed for improved efficiency, reasoning, and multilingual capabilities, offering enhanced performance for various AI applications.
  • Gemma2: The next-generation open-source model from Google, released on June 27, 2024, as an improved version of Gemma, available in 2B, 9B, and 27B parameter configurations.
  • Qwen: The series of models developed by Alibaba are powerful open-source AI models with strong multilingual understanding, containing different versions with varying parameters.
Appendix B Implementation Details
For other open-source LLMs, we use either Hugging Face weights or the official API for model inference, applying the same parameter settings as above.
Appendix C Few-Shot with Open-Source LLMs
To further evaluate the effectiveness of the few-shot learning method in improving LLMs’ ability to reconstruct vertical input content, we conduct experiments on two well-known open-source LLMs using two experimental datasets. The results are presented in Figure 9.
Figure 9: The accuracy comparison of two open-source LLMs after applying the three-shot learning method to interpret text input. The x-axis represents the model names and experimental datasets.
Our findings show that applying three-shot learning significantly enhances model accuracy. Gemma2-9B achieves an improvement of approximately 19 percentage points across both datasets, while Llama3.1-8B gains around 15 percentage points. Although these improvements are smaller than those observed in the GPT models shown in Figure 5, they still represent substantial gains, considering the base models’ initial reasoning and natural language capabilities.
Appendix D Does Explicit Prompt Work?
In real-world scenarios, natural system prompts do not explicitly instruct large language models to recognize and reconstruct vertical content in the input before generating the final output. To address this, we aim to examine how explicitly instructing the model to consider vertical content impacts its output.
In this section, we conduct an experiment by informing LLMs of the presence of vertical content in the input. Figure 11 shows an example of an explicit instruction for the sentiment classification task.
Figure 10: The accuracy comparison of two GPT models under three different prompt settings, with the accuracy after using an explicit prompt shown in green. The x-axis represents the model names and experimental datasets.
Explicit Prompt I want you to determine whether the sentiment of a sentence is positive or negative. However, some of the words in the sentence are inputted vertically into the model. I expect you to identify these words, reconstruct the original sentence, and provide the correct sentiment classification.
Figure 11: An example of an explicit instruction prompt for the sentiment classification task with vertical words input.
We conduct the experiment using two GPT models: GPT-3.5 and GPT-4. We compare the performance of explicit prompts with naive and few-shot prompts to investigate the actual impact of explicit prompts in helping LLMs recognize vertical input. The results are shown in Figure 10.
The findings indicate that although explicit prompts help LLMs mitigate the negative impact caused by vertical input to some extent, the improvement is quite small compared to the few-shot prompt. On the SST-2 dataset, GPT-3.5 improves its accuracy score by only 2 percentage points with an explicit prompt, whereas it gains 15 percentage points with a few-shot prompt. Similarly, for the GPT-4 model on the Rotten Tomatoes dataset, the improvement is only 8 percentage points with an explicit prompt, compared to 26 with a few-shot prompt.
On the one hand, although explicit prompts improve model accuracy, the improvements are not substantial. On the other hand, this type of prompt is not widely applicable to real-world scenarios, as LLMs are exposed to numerous attacks and vulnerabilities.
Input Format Dataset Accuracy(%) Original SST-2 96.67 QNLI 92.00 Rotten Tomatoes 97.33 Jigsaw Toxicity 95.00 Vertical SST-2 96.33 QNLI 92.00 Rotten Tomatoes 97.33 Jigsaw Toxicity 95.33
Table 5: The average accuracy of three human evaluations across four datasets, comparing results before and after vertical word input.
Appendix E Human vs. LLMs
To better understand the vulnerabilities of LLMs to vertical input compared to the human brain, we recruit three graduate-level students who are native English speakers to conduct experiments on four different datasets. Each student is presented with choices and required to select one, repeating this process twice—once before and once after the vertical input. Each time, the texts are shuffled to prevent memorization.
The experimental results of human classification are presented in Table 5, with accuracy computed as the average performance of the three students for each task. The results indicate that vertical input poses no challenge to human cognition, as evidenced by a minimal accuracy difference of just 0.33% across the two datasets — likely due to normal randomness rather than the input format. This demonstrates that while vertical input is easily comprehensible for humans, it remains a difficult problem for current LLMs, highlighting both the potential risks associated with relevant applications and the gap in format understanding between humans and LLMs.